OpenStudy (anonymous):

How many gallons are in 1.00 cubic feet? use dimensional analysis.

OpenStudy (anonymous):

What unit conversion factors do you have available?

OpenStudy (anonymous):

\[1cm ^{3}=1 ml\] 2.54 cm = 1 in 1.06 l = 1 qt and all the other normal conversion rates...

OpenStudy (anonymous):

Ok, then I think the path will be from feet to inches to centimeters to milliliters to liters to quarts to gallons.

OpenStudy (anonymous):

how would you convert cm3 to in?

OpenStudy (anonymous):

would you have to cube 2.54?

OpenStudy (anonymous):

Yes, and the 12 inches per foot. \[1.00ft^3 \cdot \frac{12^3inch^3}{ft^3} \cdot \frac{2.54^3cm^3}{inch^3} \cdot \frac{1ml}{1cm^3} \cdot \frac{1liter}{1000ml} \cdot \frac{1quart}{1.06liters} \cdot \frac{1gal.}{4quarts}\]

OpenStudy (anonymous):

All the units cancel except for gallons, so multiply and divide as needed and round to 3 sig.figs.

OpenStudy (anonymous):

okay. thanks...lemme calculate it and ill tell you if it works out :)

OpenStudy (anonymous):

i got 6.68...sounds okay, right

OpenStudy (anonymous):

Sounds reasonable.

OpenStudy (anonymous):

Ok, I checked it on my calculator and got the same. Do you understand how the process works now? It's all a matter of arranging the conversion factors so that the units you don't want cancel out, and the unit you're changing to is left in the numerator.